# Consumption Partial Insurance

### v5

@@ A.2 Taylor approximation to lifetime budget constraint
${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)-{\operatorname*{\mathbb{E}}}_{t-2}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)$ $\approx(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}$
$+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}$
@@
$+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.$
The linear terms in the first line are the first-order approximation of BPP (adapted for biennial data); the quadratic terms in the next lines are the refinement from the second-order approximation.
@@
where $\vartheta_{it+s}^{C}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+s}-s\ln(1+r)\right)/\sum_{\kappa=0}^{T-t}\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+\kappa}-\kappa\ln(1+r)\right)$ is an annuitization factor equal to the expected share of time $t+s$ consumption in total lifetime consumption. $\vartheta_{it+s}^{C}$ pertains to expectations at $t-2$, so it is known at $t-2$ and later.
@@
We derive this by noting that future expectations errors are zero by assumption, and that $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})=\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}\delta_{ss}+\sum_{s=0}^{T-t}\sum_{\ell\neq s}\vartheta_{it+s}^{C}\delta_{s\ell}-(\sum_{s=0}^{T-t}\vartheta_{it+s}^{C})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{C})=0$. Plugging these results in the approximation above and taking a first difference in expectations between $t-2$ and $t$ yields
@@ A.3 Biennial consumption growth
Given the biennial nature of our data, it follows from (A.1) that observed consumption growth is $\Delta^{2}c_{it+s}=\xi_{it+s}+\xi_{it+s-1}+\theta_{it+s-1}\epsilon_{it+s}+\theta_{it+s-2}\epsilon_{it+s-1}$. The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a difference between $t$ and $t-2$. Dropping the index $s$ and replacing the expectations errors with the quadratic expression in income shocks (the budget must balance so we bring the two sides of the budget constraint together), we obtain the final expression for consumption growth with biennial survey data, given by
$\begin{split}\Delta^{2}c_{it}&\approx\xi_{it}+\xi_{it-1}+(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}\\ &+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}\\ &+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.\end{split}$ (A.3)
The first line of (A.3) corresponds to the linear consumption function (4), adapted here for the biennial nature of the data and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}.$
The transmission parameters of income shocks are given by $\phi_{it}^{(1)}=1-\pi_{it}$ and $\psi_{it}^{(1)}=(1-\pi_{it})\vartheta_{it}^{Y}$, where $\pi_{it}$ and $\vartheta_{it}^{Y}$ were defined previously.
The quadratic terms in the next two lines of (A.3) reflect the refinement from the higher-order approximation. This higher-order approximation leads to the quadratic consumption function (5), adapted here for the biennial nature of the data and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\left(\zeta_{it}+\zeta_{it-1}\right)^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\left(\zeta_{it}+\zeta_{it-1}\right)v_{it}.$
The additional transmission parameters of income shocks are given by $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\pi_{it}$, $\psi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})$, and $\omega_{it}^{(22)}=(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}$.
@@ Appendix B Identification details
This appendix provides detailed identification statements based on our use of the recent biennial PSID data. There are two points of departure from section 3 in the paper.
First, we recast the income and consumption processes to reflect the biennial frequency of the modern PSID. Unexplained income growth is thus given by $\Delta^{2}y_{it}=\Delta y_{it}+\Delta y_{it-1}=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}$, with $\Delta y_{it}$ given by (3). The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a first difference in variable $x$ between $t$ and $t-2$, i.e., the observed frequency.
Second, we allow for classical measurement error. We let observed income $y_{it}^{*}$ and consumption $c_{it}^{*}$ be the sum of their corresponding true value and measurement error, namely
@@
The moments of income measurement error are not separately identifiable from the moments of the transitory shock. This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
${\operatorname*{\mathbb{E}}}((u^{y}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{y}_{t}}^{2}&\text{for }m=2\\ 0&\text{for }m=3\\ 3(\sigma_{u^{y}_{t}}^{2})^{2}&\text{for }m=4\\ \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}((u^{c}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{c}_{t}}^{2}&\text{for }m=2\\ \gamma_{u^{c}_{t}}&\text{for }m=3\\ \kappa_{u^{c}_{t}}&\text{for }m=4.\\ \end{cases}$
We retrieve the variance of income error from the validation study of the PSID in Bound et al. (1994), as we explain in section 4. We do not need similar assumptions for the consumption error, whose moments we identify below. We further assume that income and consumption errors are mutually independent (but note discussion of consumption imputation from the CEX), independent over time, and independent of the income shocks and taste heterogeneity.
These choices lead to our processes for observed income growth, consumption growth in the linear specification, and consumption growth in the quadratic specification, given by
$\Delta^{2}y_{it}$ $=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}+\Delta^{2}u^{y}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}+\Delta^{2}u^{c}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}$
$+\phi_{it}^{(2)}(\zeta_{it}+\zeta_{it-1})^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}(\zeta_{it}+\zeta_{it-1})v_{it}+\Delta^{2}u^{c}_{it},$
respectively. These are the theoretical expressions we bring to the data.
Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks within a two-year period. We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distributions are stationary. Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$. For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, and $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$. These last restrictions are not needed for identification with biennial data, but they simplify the illustration.
@@ B.1 Income process parameters
$\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
$\gamma_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\gamma_{v_{t}}$ $=-{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it+2}).$
@@
$\kappa_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})-\kappa_{v_{t}}-3(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
$-12\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-12\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it+2})^{2})-4(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
$-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-8\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}.$

### v6

@@ A.2 Taylor approximation to lifetime budget constraint
${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)-{\operatorname*{\mathbb{E}}}_{t-2}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)$ $\approx(1-\pi_{it})\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})$
$+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}$
$+\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}$
@@
$+(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}.$
The linear terms in the first two lines are the first-order approximation of BPP (adapted here for biennial data and income persistence); the quadratic terms in the next lines are the refinement from the second-order approximation.
@@
where $\vartheta_{it+s}^{C}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+s}-s\ln(1+r)\right)/\sum_{\kappa=0}^{T-t}\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+\kappa}-\kappa\ln(1+r)\right)$ is an annuitization factor equal to the expected share of time $t+s$ consumption in total lifetime consumption. $\vartheta_{it+s}^{C}$ pertains to expectations at t2t-2, so it is known at t2t-2 and later.
@@
We derive this by noting that future expectations errors are zero by assumption, and that $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})=\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}\delta_{ss}+\sum_{s=0}^{T-t}\sum_{\ell\neq s}\vartheta_{it+s}^{C}\delta_{s\ell}-(\sum_{s=0}^{T-t}\vartheta_{it+s}^{C})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{C})=0$. Plugging these results in the approximation above and taking a first difference in expectations between t2t-2 and $t$ yields
@@ A.3 Biennial consumption growth
Given the biennial nature of our data, it follows from (A.1) that observed consumption growth is $\Delta^{2}c_{it+s}=\xi_{it+s}+\xi_{it+s-1}+\theta_{it+s-1}\epsilon_{it+s}+\theta_{it+s-2}\epsilon_{it+s-1}$. The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a difference between $t$ and t2t-2. Dropping the index $s$ and replacing the expectations errors with the quadratic expression in income shocks (the budget must balance so we bring the two sides of the budget constraint together), we obtain the final expression for consumption growth with biennial survey data and income persistence, given by
$\begin{split}\Delta^{2}c_{it}&\approx\xi_{it}+\xi_{it-1}+(1-\pi_{it})\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}\\ &+\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}\\ &+(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}.\end{split}$ (A.3)
The first line of (A.3) corresponds to the linear consumption function, adapted here for income persistence and the biennial nature of the data, and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}.$
The transmission parameters of income shocks are given by $\phi_{it}^{(1)}=(1-\pi_{it})\Theta_{it}^{Y}$ and $\psi_{it}^{(1)}=(1-\pi_{it})\vartheta_{it}^{Y}$. In the baseline, the permanent-transitory income process has $\rho=1$ and further restricts $\phi_{it}^{(1)}=1-\pi_{it}$ because $\Theta_{it}^{Y}=\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$.
The quadratic terms in the next two lines of (A.3) reflect the refinement from the higher-order approximation. This leads to the quadratic consumption function, adapted here for income persistence and the biennial nature of the data, and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}\\ +\phi_{it}^{(2)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)v_{it}.$
The additional transmission parameters of income shocks are given by $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}$, $\psi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})$, and $\omega_{it}^{(22)}=(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}$. The baseline permanent-transitory income process has $\rho=1$ and further restricts $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\pi_{it}$ and $\omega_{it}^{(22)}=(1-\pi_{it})\pi_{it}\vartheta_{it}^{Y}$, because $\widetilde{\Theta}_{it}^{Y}=\Theta_{it}^{Y}=\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$.
@@ Appendix B Identification details
This appendix provides detailed identification statements based on our use of the recent biennial PSID data. There are three points of departure from section 3 in the paper.
First, we use the most general AR(1) specification for income, shown in section 6. The baseline permanent-transitory specification is a special case with $\rho=1$. Second, we allow for classical measurement error. We let observed income $y_{it}^{*}$ and consumption $c_{it}^{*}$ be the sum of their corresponding true value and measurement error, namely
@@
The moments of income measurement error are not separately identifiable from the moments of the transitory shock. This is why we assume the income error to be Gaussian (but let the consumption error unrestricted), so we specify
[…]
We retrieve the variance of income measurement error from the validation study of the PSID in Bound et al. (1994), as we explain in section 4. We do not need similar assumptions for the consumption measurement error, whose moments we identify below. We further assume that income and consumption errors are mutually independent (but note the discussion of consumption imputation from the CEX in appendix C.4), independent over time, and independent of the income shocks and taste heterogeneity. Third, we recast the income and consumption processes to reflect the biennial frequency of the modern PSID.
These choices lead to the theoretical expressions that we bring to the data, namely the process for income growth, the quasi-difference in income, consumption growth in the linear specification, and consumption growth in the quadratic specification, given respectively by
$\Delta^{2}y_{it}$ $=(\rho^{2}-1)P_{it-2}$~ $+$ $~\zeta_{it}+\rho\zeta_{it-1}+\Delta^{2}v_{it}+\Delta^{2}u^{y}_{it},$
$\Delta^{2}_{\rho}y_{it}$ $=$ $~\zeta_{it}+\rho\zeta_{it-1}+\Delta^{2}_{\rho}v_{it}+\Delta^{2}_{\rho}u^{y}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}$ $+$ $~\phi_{it}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})+\psi_{it}^{(1)}v_{it}+\Delta^{2}u^{c}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}$ $+$ $~\phi_{it}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})+\psi_{it}^{(1)}v_{it}$
$+$ $~\phi_{it}^{(2)}(\zeta_{it}+\rho\zeta_{it-1})^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}(\zeta_{it}+\rho\zeta_{it-1})v_{it}+\Delta^{2}u^{c}_{it}.$
The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a first difference in $x$ between $t$ and t2t-2, i.e., the observed frequency. The quasi-difference is defined as $\Delta^{2}_{\rho}x_{t}=\Delta_{\rho}x_{t}+\rho\Delta_{\rho}x_{t-1}=x_{t}-\rho^{2}x_{t-2}$, where $\Delta_{\rho}x_{t}=x_{t}-\rho x_{t-1}$ is the quasi-difference between $t$ and t1t-1.
Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks within a two-year period. We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distributions are stationary. Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$. For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$, and $\kappa_{v_{t}}\approx\kappa_{v_{t-2}}$. These last restrictions are not needed for identification, but they simplify the illustration.
@@ B.1 Income process parameters
$\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{1+\rho^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-\frac{1}{\rho^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
$\gamma_{\zeta_{t}}$ $=\frac{1}{1+\rho^{3}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa})\qquad$ $\text{and}\qquad\gamma_{v_{t}}$ $=-\frac{1}{\rho^{2}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it+2}).$
@@
$\kappa_{\zeta_{t}}$ $=\frac{1}{1+\rho^{4}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})-\frac{1+\rho^{8}}{1+\rho^{4}}\kappa_{v_{t}}-6\frac{\rho^{2}}{1+\rho^{4}}(\sigma_{\zeta_{t}}^{2})^{2}-6\frac{\rho^{4}}{1+\rho^{4}}(\sigma_{v_{t}}^{2})^{2}-3(1+\rho^{4})(\sigma_{u^{y}_{t}}^{2})^{2}$
$-6(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-6(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-6(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$
$\kappa_{v_{t}}$ $=\frac{1}{\rho^{4}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it+2})^{2})-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(\sigma_{\zeta_{t}}^{2})^{2}-\frac{1+\rho^{4}+\rho^{8}}{\rho^{4}}(\sigma_{v_{t}}^{2})^{2}-\frac{1+4\rho^{4}+\rho^{8}}{\rho^{4}}(\sigma_{u^{y}_{t}}^{2})^{2}$
$-2(1+\rho^{2})\frac{1+\rho^{4}}{\rho^{4}}\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-2(1+\rho^{2})\frac{1+\rho^{4}}{\rho^{4}}\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-\left(2\frac{(1+\rho^{4})^{2}}{\rho^{4}}+4\right)\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}.$
In all cases, either we need an external estimate for $\rho$ (e.g., $\rho=1$ as in the permanent-transitory process) or treat $\rho$ as a free parameter to be jointly estimated with the moments of income shocks. In the latter case, $\rho$ is pinned down by
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})={\operatorname*{\mathbb{E}}}((y_{it}-\rho^{2}y_{it-2})\times(y_{it-4}-y_{it-6}))=0,$
namely a “long” intertemporal covariance. Note that when $\rho$ is a free parameter, five time period of data are needed at a minimum (i.e., t6t-6 through $t+2$ biennially) to identify both the variance of persistent shocks and $\rho$ itself, versus four periods when $\rho$ is set externally.